A random sample of size 20 is taken, resulting in a sample mean of 16.45 and

a sample standard deviation of 3.59. Assume x is normally distributed and use

this information and α=0.5 to test the following hypotheses. (use critical

t.025,19 = ±2.093)


Ho: μ = 16

Ha: μ 16


1
Expert's answer
2021-08-02T12:56:08-0400

In this case,

"Sample\:mean=\overline{x}=16.45"


"Sample\:standard\:deviation=s=3.59\:"


"Sample\:size=n=20"


"\:Degrees\:of\:freedom=19"


"Significance\:level=\alpha =0.05"


Step 1: "H_0:\mu=16"

"H_1:\mu\:\:\not= \:16" "(Two" "tailed" "test)"

Step 2: "The\:test\:statistic\:is,\:t=\frac{\overline{x}-\mu }{\left(\frac{s}{\sqrt{n}}\right)}\:"

"t=\frac{16.45-16}{\left(\frac{3.59}{\sqrt{20}}\right)}=0.5606"

Step 3: "P-value=P\left(\left|t\right|>0.5606\right)=0.5816"

Step 4: Conclusion: Since the "P-value" is greater than the significance level, we fail to reject "H_o" .